A train travels at a rate of (4x+5) miles per hour. How many miles can it travel at that rate in (x-1) hours?
i have tried this question multiple times and eneded up getting the same 4 answers:
3x-4 miles
5x-4 miles
4x(squared)+x-5 miles
and 4x(squared)-9x-5 miles

Answer by Earlsdon(6287)

Use: d = r*t where d = distance (miles), r = rate (mph), and t = time of travel (hrs).
d = r*t Substitute r = (4x+5)mph and t = (x-1)hrs.
d = (4x+5)*(x-1) Use FOIL to multiply.
d = 4x^2-4x+5x-5 Simplify.
d = 4x^2+x-5 miles.